# Scale-dependent chirality as a smoking gun for Abelian gauge fields during inflation

### v1

@@ 5 Discussion and Conclusions
In this work we focused on two-point function and so one might wonder how would our findings reflect on higher-order correlators. This question is relevant already at the level of currently operating experimental probes as existing bounds on GW non-Gaussianities from the Planck mission have placed constraints on axion gauge field models [100]. Refining purely GW [65] as well as mixed non-Gaussianity [57] calculations by means of our semi-analytical or fully numerical approach will be certainly worthwhile. As repeatedly mention in the text, another natural arena to revisit is that of PBH production and the corresponding scalar-induced gravitational wave spectrum [54, 55, 84, 60]. We leave this to upcoming work.
@@ Appendix B Scalar and tensor perturbations induced by the vector fields
where we utilized the following identity
$\left|\epsilon^{(-)}_{i}(\mathbf{q})\epsilon^{(-)}_{i}(\mathbf{q}^{\prime})\right|^{2}=\frac{1}{4}\left(1-\hat{\mathbf{q}}.\hat{\mathbf{q}}^{\prime}\right)^{2}.$ (B.19)
Using (B.18), we then take the two point function of the sourced curvature perturbation (B.16) following the same steps we carried for the tensor perturbations. In this way, the sourced power spectrum of curvature fluctuations takes the following form
@@
We note that to derive this expression, we only took the late time limit of the outer most Green’s function in (B.16) using (B.3) while we kept the full expression (B.2) for the inner Green’s function. Following the discussion we provide in Appendix A, to fully numerically compute (B.21), we also require
@@
For the evaluation of the sourced curvature power spectrum in the WKB approximation, we use the expression use (B) and (B.24) with the numerically obtained normalization factors $N(\xi_{*},x_{*},\delta)$ we describe in the previous section to discuss phenomenological implications of the resulting signal (see section 4).

### v2

@@ 5 Discussion and Conclusions
In this work we focused on two-point function and so one might wonder how would our findings reflect on higher-order correlators. This question is relevant already at the level of currently operating experimental probes as existing bounds on GW non-Gaussianities from the Planck mission have placed constraints on axion gauge field models [117]. Refining purely GW [80] as well as mixed non-Gaussianity [72] calculations by means of our semi-analytical or fully numerical approach will be certainly worthwhile. As repeatedly mention in the text, another natural arena to revisit is that of PBH production and the corresponding scalar-induced GW spectrum [69, 70, 100, 75]. We leave this to upcoming work.
@@ Appendix B Scalar and tensor perturbations induced by the vector fields
Dominant contribution to the integral and UV cut-off insensitivity.
We devote this sub-section to dive deeper into some important features of the above-mentioned correlators. For illustrative purposes, we will use Fig. 11 as a typical profile of the particle production parameter $\xi$ considered in this work. Let us denote by $k_{A}$ the momentum of an internal gauge field perturbation which leaves the horizon at the moment $N_{A}$ and likewise for other time points shown in Fig. 11: $k_{*}$ at $N_{*}$ and $k_{B}$ at $N_{B}$, where $N_{A}<N_{*}<N_{B}$. Next, let us consider an external gravitational wave of momentum $k_{A}$ which, following our definitions above, leaves the horizon at moment $N_{A}$. In principle, such a GW can be sourced by gauge fields of all possible (internal) momenta $p$, from very small (those that left the horizon much earlier than $N_{A}$) to very large momenta (those that leave the horizon after $N_{B}$). The conventional wisdom so far has been that the dominant contribution to a sourced GW of external momentum $k_{A}$ comes from gauge fields of internal momenta in the neighborhood of $p\sim k_{A}$ since much larger or smaller momenta will suffer from significant phase-space suppression.
If that were the case in the present context, one could rightfully point out that our cut-off procedure would not be reliable in light of recent works such as [44]. Therein, it has been conclusively shown that, for small particle production parameters (as is the case at $N_{A}$), the simple cut-off procedure is ambiguous and one ought to use a more robust regularization and renormalization prescription to extract the final result as done in [118, 119]. One of the most important results of this work is that such conventional wisdom does in fact not apply when $\xi$ displays a bumpy profile as in Fig. 11. Instead, in such cases all sourced GWs, regardless of their momenta (and correspondingly horizon-crossing time), and by extension the value of the particle production parameter $\xi$ in that moment, are universally sourced by gauge fields of momentum in the neighborhood of $k_{*}$ which crossed the horizon at around moment $N_{*}$. It follows that all GWs are correspondingly enhanced by the maximum $\xi$ value, i.e. $\xi_{*}$ , which is sufficiently large to make our cut-off procedure reliable.
Figure 11: The typical time-evolution of the “bumpy” particle production parameter in the cases considered in this work. The labels $N_{A}$, $N_{*}$ and $N_{B}$ denote three distinct moments in time: one for which $\xi$ has not yet reached its maximum value, one for which $\xi$ is exactly at the maximum $\xi_{*}$, and one for which $\xi_{*}$ is decreasing from the maximum. The horizontal axis denotes the number of e-folds relative to $N_{*}\,$.
One can see that this is the case by considering that the sourced GWs are ultimately integrals over four powers of the gauge mode functions. Parametrically then, the amplitude of the sourced GW is proportional to $P_{h}\propto{\rm PSS}\times{\rm e}^{4\pi\xi}$, where PSS stands for Phase Space Suppression. Since gauge fields of all possible momenta source the GWs, there is a trade-off between the size of the particle production parameter and the corresponding phase space suppression due to the hierarchy between the external (GW) and internal (gauge field) momentum. Depending on the width and the height of the peak in the $\xi$ profile, it is possible to identify an example for which the phase space suppression is the subdominant effect. It is instead the enormous occupation number attained by the gauge field at moment $N_{*}$ that dominates the production of GWs of all scales. Pointing out that such cases exists and they respect the adiabatic-evolution conditions for $\xi$ is one of the main points of the present work.
As a side note, we find it important to stress that the GWs of (e.g.) momentum $k_{A}$ are still sourced by gauge fields of momentum $p\sim k_{A}$, for which there is an ambiguity regrading the regularization-renormalization procedure. However, any such contribution coming from small gauge field momenta is a negligible part of the total result. We therefore do not expect our results to significantly change upon implementing a more sophisticated approach than the simple cut-off we employ here.
As a final note, and also a segue to the next sub-section, it is worthwhile to note that (i) GWs of momentum $k_{A}$ are predominantly sourced by gauge field momenta that are much greater than $k_{A}$, (ii) GWs of momentum $k_{*}$ are predominantly sourced by gauge fields in the neighborhood of $k_{*}$, and finally (iii) GWs of momentum $k_{B}$ are predominantly sourced by at least one gauge field whose momentum is far smaller than $k_{B}$. We will see that this hierarchy between the size of the external momentum $k$ and the loop momenta $p$ (as well as their angles) is precisely the physical origin of the scale-dependent chirality which we highlight in this work.
Revisiting the origin of scale-dependent GW chirality. We find it useful to provide here a rather condensed discussion on the emergence of scale-dependent parity violation of the produced GW signal. More details have been provided in section 4.2. For this purpose, we turn our attention to the part of the two-point GW correlator called $f_{2,\lambda}$ that exhibits scale-dependence. The variable $x_{*}=k/k_{*}$, in turn, parametrizes the size of the external tensor modes in eq. (B.13). In this expression we can identify three important ingredients that can help us gain a better understanding of the origin of scale-dependent GW chirality:
  1. 1.
    The dominant support to the tensor correlator (B.13) comes from the peak of the scale-dependent vector normalization factors $N$ (at fixed $\xi_{*}$ and $\delta$), which depend on the normalized loop momenta $\tilde{p}$, the cosine angle $\eta$ and the wave-number $x_{*}=k/k_{*}$ of the external GW.
  2. 2.
    The part of the tensor correlator that determines the amplitude of the different GW polarization states is given by
    $p_{\lambda}(\tilde{p},\eta)\equiv\left(1-\lambda\eta\right)^{2}\left(1-\lambda\frac{(1-\tilde{p}\eta)}{\sqrt{1-2\tilde{p}\eta+\tilde{p^{2}}}}\right)^{2}\;.$ (B.16)
  3. 3.
    A tensor mode with momentum $\vec{k}$ is produced via a loop of two vector modes with momenta $\vec{k}-\vec{p}$ and $\vec{p}$, such that the momentum is conserved at each vertex.
We infer from the first point above that to produce a soft GW mode with $x_{*}=k/k_{*}\ll 1$, the vector modes running in the loop must satisfy $\tilde{p}\gg 1$. This becomes apparent from the expression (B.13) by noticing the fact the integral gets most of its support from the peak of the vector normalization factor $N$ where $\xi_{*}$ is at its maximum (see Figure 2). In this limit ($\tilde{p}\to\infty$), momentum conservation (see point 3 above) at the vertex singles out the anti-aligned configuration [^15] Notice that such a configuration can be achieved for any angle $\eta$ since $\eta$ is the angle between the external $\vec{k}$ and internal momenta $\vec{p}$. between the two internal gauge modes. The key point is that this configuration is not sensitive to left/right handed GW modes as each carries equal weight (see point 2 above):
$\lim_{\tilde{p}\to\infty}\frac{p_{-}(\tilde{p},\eta)}{p_{+}(\tilde{p},\eta)}\quad\to\quad 1+\frac{4}{\tilde{p}}\quad\to\quad 1\;\;.$ (B.17)
In other words, in the anti-aligned configuration the ratio of the phase space available for the two GW polarization states approaches unity, $p_{+}/p_{-}\to 1$, leading to a suppressed chirality in the IR, $x_{*}=k/k_{*}\ll 1$.
Similarly, the GW spectrum (B.13) in the UV ($x_{*}=k/k_{*}\gg 1$) only has support for one of the two loop momenta satisfying $\tilde{p}\ll 1$ as the vector sources are maximal only for $\tilde{p}\,x_{*}\sim 1$. In this limit ($\tilde{p}\to 0$), momentum conservation allows for any angle between the internal gauge modes. However, the co-linear configuration allows only one polarization, forbidding altogether the production of the opposite one. Taking the limit again we find
$\lim_{\tilde{p}\to 0}\frac{p_{-}(\tilde{p},\eta)}{p_{+}(\tilde{p},\eta)}\quad\to\quad\frac{16}{(\eta-1)^{4}\tilde{p}^{4}}\quad\to\quad\infty\;\;,$ (B.18)
which confirms our findings: GW spectra exhibit maximal chirality for scales following the peak, i.e. scales such that $k/k_{*}\gg 1$ (see Figs. 7 and 8).
Figure 12: A plot of the ratio of the two polarization-dependent factors defined in (B.16). The greater the hierarchy between internal and external momentum ($\tilde{p}\gg 1$), the more the sourcing of the GWs becomes chirality-insensitive. The red dots mark the points in the phase space that dominate the sourcing of GWs that left the horizon in the three moments marked in Fig. 11.
In order to emphasise our results, we plot the ratio of the contributions from the two different polarizations in Fig. 12, a representation fully consistent with the limits evaluated analytically above. The final GW power will depend on the product between the polarization dependent factor defined in (B.16) and the polarization independent residual factor in the integrand of eq. (B.11). This residual factor is precisely what controls which part of the phase space $(\tilde{p},\eta)$ will dominate the GW sourcing. We identified the point in phase space at which the production peaks for the three sample moments in time in the toy model of Fig. 11. We indicate these points with red dots in Fig. 12 to show more explicitly how the hierarchy between the external and internal momenta have a direct impact on the GW chirality. Indeed, based on the locations of peak production, one would expect the chirality of the sourced GWs to increase as one moves from $N_{A}$ to $N_{*}$ and finally to $N_{B}$. Such a behaviour is reflected in the results we report in the main text.
@@
where we utilized the following identity $\left|\epsilon^{(-)}_{i}(\mathbf{q})\epsilon^{(-)}_{i}(\mathbf{q}^{\prime})\right|^{2}=\left(1-\hat{\mathbf{q}}.\hat{\mathbf{q}}^{\prime}\right)^{2}/4$. Using (B.21), we then take the two point function of the sourced curvature perturbation (B.19) following the same steps we carried for the tensor perturbations. In this way, the sourced power spectrum of curvature fluctuations takes the following form
@@
Note that here, we only take the late time limit of the outer most Green’s function in (B.19) using (B.3) while keeping the full expression (B.2) for the inner Green’s function. Following the discussion we provide in Appendix A, to fully numerically compute (B.23), we also require
@@
To evaluate the sourced power spectrum within WKB method, we use (B) and (B.26) with the numerically obtained normalization factors $N(\xi_{*},x_{*},\delta)$ we describe in the previous section to discuss phenomenological implications of the resulting signal (see section 4).